OpenStudy (anonymous):

A solution is made by dissolving 21.5 grams of glucose (C6H12O6) in 255 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 °C/m? Show all of the work needed to solve this problem.

OpenStudy (anonymous):

Change in freezing point=Kf * m. m- molality of glucose. m= no. of moles of glucose/mass of solvent(water) in kg. no. of moles of glucose = 21.5/180.15= 0.1193 m= 0.1193/0.255 = 0.468 moles/kg. Change in freezing point = (-1.86)*(0.468) = -0.8705 °C Actually the freezing point increases by 0.8705 °C .